# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v1

@@ III-C Comparing REINFORCE with PPO
$=\mathbb{E}_{\tau\sim p_{\mathbf{\theta}}(\tau)}\left[\sum_{t=0}^{T}\nabla_{\mathbf{\theta}}\log\pi_{\mathbf{\theta}}R(\tau)\right],$ (5)
@@
where $\tau$ represents a trajectory, $p_{\mathbf{\theta}}(\tau)$ represents the probability of trajectory $\tau$ occurring under policy $\pi_{\mathbf{\theta}}$, and $R(\tau)$ represents the cumulative reward of trajectory $\tau$. Each state in the MDPs for formulaic alpha factor mining processes corresponds to a sequence of tokens denoting the currently generated part of the expression. The initial state is always BEG, so a valid state always starts with BEG and is followed by previously chosen tokens, which means $\mathbf{s}_{t}=\mathbf{a}_{1:t-1}$, the current policy is $\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$, the score function $s_{\mathbf{\theta}}\left(\mathbf{a}_{1:t}\right)=\nabla_{\mathbf{\theta}}\log\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$, and $p_{\mathbf{\theta}}(\tau)=p\left(a_{0}\right)\prod_{t=0}^{T}\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$. Since IC can only be calculated after the complete expression is generated, only the reward for the final step is non-zero. Therefore, $R(\tau)=r\left(\mathbf{a}_{1:T}\right)$.The Policy Gradient (5) can then be rewritten as:
$g(\mathbf{\theta})$ $=\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\mathbf{\theta}}}\left[\sum_{t=0}^{T}\nabla_{\mathbf{\theta}}\log\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)r\left(\mathbf{a}_{1:T}\right)\right]$
$=\mathbb{E}_{\mathbf{a}_{1:t}\sim\pi_{\mathbf{\theta}}}\left[\sum_{t=0}^{T}s_{\mathbf{\theta}}\left(\mathbf{a}_{1:t}\right)r\left(\mathbf{a}_{1:T}\right)\right].$ (6)
@@
$r\left(\mathbf{a}_{1:t}\right)$ is used as the metric of the effectiveness of the current policy. It can also be replaced by the following five metrics: $\sum_{t=0}^{\infty}r_{t}$ (total reward of the trajectory)[45], $\sum_{t^{\prime}=t}^{\infty}r_{t^{\prime}}$ (reward following action $a_{t}$)[45], $\sum_{t^{\prime}=t}^{\infty}r_{t^{\prime}}-b(\mathbf{s}_{t})$ (baselined version of the previous formula)[45], $Q^{\pi}(\mathbf{s}_{t},a_{t})$ (state-action value function)[46], $r_{t}+V^{\pi}(\mathbf{s}_{t+1})-V^{\pi}(\mathbf{s}_{t})$ (TD residual)[47], and $A\left(\mathbf{s}_{t},a_{t}\right)=Q^{\pi}(\mathbf{s}_{t},a_{t})-V^{\pi}(\mathbf{s}_{t})$ (advantage function)[47]. Advantage function is used to measure the effectiveness of the policy in PPO, which implies that the TD algorithm[48] can be used to estimate the policy gradient. The advantage of using TD instead of Monte Carlo algorithms[49] to update network weights is that data from each step can be used for training, increasing sample efficiency. Additionally, to enable the reuse of the samples for learning, as in off-policy algorithms, PPO employs importance sampling:
$L_{surr}\left(\mathbf{\theta}\right)=\mathbb{E}_{\mathbf{a}_{1:t}\sim\pi_{\mathbf{\theta}_{\text{old }}}}\left[\sum_{t=1}^{T}A\left(\mathbf{a}_{1:t}\right)\psi\left(\mathbf{a}_{1:t}\right)\right],$ (7)
where ratio $\psi\left(\mathbf{a}_{1:t}\right)=\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)/\pi_{\mathbf{\theta}_{\text{old }}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$ is the importance weight. Although $\pi_{\mathbf{\theta}_{\text{old }}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$ can be any distribution, the expectation of two random variables being the same does not imply that their variances are also the same. When the difference between $\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$ and $\pi_{\mathbf{\theta}_{\text{old }}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$ is too large, it can introduce significant variance. Therefore, the PPO algorithm incorporates a clipping operation based on the policy gradient theorem and importance sampling to ensure that the difference between the two distributions is not too large, resulting in the final PPO loss function:
$L_{surr}\left(\mathbf{\theta}\right)=\mathbb{E}_{\mathbf{a}_{1:t}\sim\pi_{\mathbf{\theta}_{\text{old}}}}\left.\Bigg{[}\sum_{t=1}^{T}A\left(\mathbf{a}_{1:t}\right)\min\left.\Big{(}\psi\left(\mathbf{a}_{1:t}\right),\right.\right.$
$\left.\left.\operatorname{clip}\left(\psi\left(\mathbf{a}_{1:t}\right),1-\delta,1+\delta\right)\right.\Big{)}\right.\Bigg{]}.$ (8)
However, for these MDPs where non-zero rewards are only received at the final step, algorithms such as REINFORCE, as well as our proposed QFR algorithm, are more appropriately suited. We will delve into the details in Section IV-C. Before that, let us briefly introduce the respected REINFORCE algorithm. Based on the previously mentioned policy gradient (6), an unbiased estimate of this gradient can be obtained using the Monte Carlo method:
@@
where $a_{t}^{i}\sim\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{a}_{1:t-1}^{i}\right)$ for all $t=1,\ldots,T$. Performing gradient ascent can update $\mathbf{\theta}$ to optimize the performance of the policy network: $\mathbf{\theta}_{k+1}\leftarrow\mathbf{\theta}_{k}+\eta_{k}\widehat{g}\left(\mathbf{\theta}_{k}\right)$, where $\eta_{k}$ represents the learning rate at the $k$-th iteration, which is essentially a likelihood maximization weighted by rewards using samples extracted from rollouts.
IV Better RL in Formulaic Alpha Factors Mining
In this section, we introduce an innovative reinforcement learning algorithm for alpha factor mining, termed QuantFactor REINFORCE (QFR). We provide a theoretical analysis that establishes an upper bound on the variance of the QFR algorithm in MDPs modeling the formulaic alpha factor mining process. We demonstrate that QFR exhibits reduced variance compared to the celebrated REINFORCE algorithm. The proof that QFR, when applied to MDPs with trajectory-level rewards, exhibits the lowest variance compared to those with non-trajectory-level rewards, is also provided. Furthermore, we introduce IR as a reward shaping mechanism to promote the generation of steady alpha factors.
Figure 2: Comparison between QFR and PPO. QFR requires rollouts once with sampling settings and once with greedy settings to removes all the components related to the value model in PPO. This significantly improves the overall convergence speed.
@@ IV-A The Proposed Algorithm
Based on the MDPs defined in Section III-B for formulaic alpha factors mining, unlike the work in [14], we use Monte Carlo methods to obtain an estimate of the policy gradient (6) to optimize a policy $\pi_{\mathbf{\theta}}(a\mid\mathbf{a}_{1:t-1})$. Inspired by the REINFORCE algorithm with baseline[16], our QFR algorithm also modifies the gradient estimate by introducing a subtractive baseline value to exhibit a smaller variance:
@@
The baseline value $r\left(\mathbf{\bar{a}}_{1:T}^{i}\right)$ can be obtained by greedily sampling a formulaic alpha factor and calculating the associated reward, where $\bar{a}_{t}\in\operatorname{argmax}\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{\bar{a}}_{1:t-1}\right)$. The proposed baseline serves a normalizing function by comparing the reward of the random response with that of the greedy response. It helps reduce the variance in the gradient estimate. We will provide a detailed theoretical analysis in Section IV-C. Notably, while formally our baseline has the same form as in work [16], the mechanism by which this baseline works is quite different for the task in work [16] than it is for our MDPs here. To our best knowledge, we are the first to introduce this baseline in MDPs for formulaic alpha factors mining, and we provide novel and detailed theoretical proofs of why this baseline is particularly well suited to these MDPs in Section IV-C
Our QFR algorithm builds on the computational efficiency of REINFORCE but significantly enhances its performance. Specifically, REINFORCE often faces large variance in stochastic gradients during training. To address these challenges, QFR employs a greedy baseline value that adapts to trajectory-level rewards. While PPO can also reduce large variance during training, it does so at the cost of introducing a value model that matches the scale of the policy model. This substantially increases the sampling time, as a forward pass must be performed for each state sampled from the buffer to obtain $V^{\pi}(\mathbf{s}_{t})$. The pipeline for the QFR algorithm is shown in Fig. 3.
Figure 3: The pipeline for the QFR algorithm. The policy model generates one token at each step, and the sequence of tokens can be uniquely converted into an expression based on RPN. The policy model uses random sampling and greedy sampling to complete two trajectories and generate two expressions, i.e., two factors. The combination model is used to maintain a weighted combination of principal factors and, at the same time, to evaluate these two factors. Specifically, the combined factor values are used to calculate $r\left(\mathbf{a}_{1:T}\right)$ and $r\left(\mathbf{\bar{a}}_{1:T}\right)$ along with asset data. These rewards, together with $\log\pi_{\mathbf{\theta}}$, form the rollout used to calculate the estimate (10).
Input: Real asset price dataset $\mathcal{Y}$, Real asset feature dataset $\mathcal{X}$, Initial policy weight $\mathbf{\theta}$, Initial combination model weight $\mathbf{\omega}$, Reward balancing coefficient $\lambda$.
@@
1while not convergent do
2       Collect a set of normal expressions $\mathcal{D}_{n}$ and a set of greedy expressions $\mathcal{D}_{g}$ by running current policy $\pi_{\mathbf{\theta}}$;
3       Compute factor values by applying $\mathcal{D}_{n}$ in $\mathcal{X}$;
4       Compute greedy factor values by applying $\mathcal{D}_{g}$ in $\mathcal{X}$;
5       Compute combined factor values and combined greedy factor values using $\mathbf{\omega}$ and formal generated alpha factors;
6       Compute rewards $r\left(\mathbf{a}_{1:T}\right)$ and greedy reward $r\left(\mathbf{\bar{a}}_{1:T}\right)$ via Eq. (14);
@@ 4       Compute greedy factor values by applying $\mathcal{D}_{g}$ in $\mathcal{X}$;
After completing an episode that generates a formulaic alpha factor, we combine this factor with other factors in the factor pool and calculate the combined factor value, using the linear factor combination model from work[50]. Due to significant differences in the scales of different formulaic alpha factor values, they are first standardized to have a mean of 00 and a maximum value of $1$. The loss of the combination model is defined as the mean squared error (MSE) between the model output and the actual asset trend:
$L(\mathbf{\omega})=\frac{1}{T}\sum_{t=1}^{T}\left\|\mathbf{z}_{t}^{\prime}-\mathbf{y}_{t}\right\|^{2},$ (11)
where $\mathbf{\omega}$ is weight of combination model. After the policy model outputs a new formulaic alpha factor, this alpha is first added to the factor pool and assigned a random initial weight. Gradient descent is then performed to optimize the weights for these formulaic alphas. If the number of factors in the factor pool exceeds a certain threshold, the formulaic alpha factor with the smallest weight and its corresponding weight will be discarded.
Figure 4: Clip function values versus training time steps, with an example of $\overline{IR}$ values. When the clipped function value, i.e., the $\overline{IR}$ test value, is greater than the $\overline{IR}$ value of the generated combined factor, the factor quality is poor, and the last step will obtain a discounted $\overline{IC}$ as a reward. Conversely, a complete $\overline{IC}$ is obtained as a reward.
IV-B Reward Shaping
The reward function from the work [14] focuses on the absolute excess returns of factors, while ignoring the risk- adjusted characteristics of the returns. Unlike this work, in order to better balance returns and risks, our work not only evaluates the predictive accuracy of factors but also considers the volatility of their predictive signals, thus providing a more comprehensive assessment of factor performance.
Specifically, we introduce IR for reward shaping. IR is a metric that measures the excess returns of factors relative to their tracking error and is commonly used to evaluate the risk-adjusted performance of the factors:
@@ IV-B Reward Shaping
(12) describes how the IR of a single factor is calculated. After the combination model, the IR for the factors becomes:
@@
(14) is the shaped reward. and II denotes the indicator function, which takes the value of $1$ if the condition inside the parentheses is satisfied, and 00 otherwise. $\alpha$ is the time delay, set to 180,000 time steps, $\eta$ is the slope of the change in the IR test value, set to 1/1,000,000, $\delta$ is the maximum IR test value, set to 0.3, and $\lambda$ is used to balance the main reward with the shaped reward, set to 0.02. The clip function is illustrated in Fig. 4.
Using the IR test as reward shaping can guide the QFR algorithm to focus more on the long-term stability and consistency of the factors during optimization, rather than just the short-term profit peaks. This reward shaping method guides the policy network to find factors that maintain stable predictive capability across different market conditions, reducing the risk of overfitting the training data. It helps to reduce the model’s over-reliance on specific market conditions, thereby enhancing the generalization ability of factors across different market environments. The pseudo-code for QFR is provided in Algorithm 1 (also see Fig. 2).
@@ IV-C The Theoretical Analysis
We provide a set of theoretical results for Algorithm 1, which includes the derivation of the upper bound of the variance, the demonstration that the variance decreases relative to the REINFORCE algorithm and the analysis of the training variance under state transition functions with various distributions. To justify the algorithm design, we first prove that the variance of the proposed algorithm is bounded.
@@
where (15) follows from the so-called log-derivative trick, (16) follows from the idea of Bartlett identity $\sum_{z}\nabla_{\mathbf{\theta}}p_{\mathbf{\theta}}(z)b=\nabla_{\mathbf{\theta}}\left[\sum_{z}p_{\mathbf{\theta}}(z)b\right]=\nabla_{\mathbf{\theta}}[1\cdot b]=0$, in which $p_{\mathbf{\theta}}$ can be any distribution and $b$ is constant[51]. Notice that $\mathbf{\bar{a}}_{1:T}$ is conditionally independent on $\mathbf{\theta}$, due to the greedy sampling and apply $p_{\mathbf{\theta}}$ to the the distribution $\pi_{\mathbf{\theta}}\left(\bar{a}_{1:T}\mid x\right)$. Then, regarding the second item, we only need to consider $\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\theta}}\left[\nabla_{\theta}\sum_{t=1}^{T}\log\pi_{\theta}(a_{t}\mid\mathbf{a}_{1:t-1})\right]=\nabla_{\theta}\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\theta}}\left[\sum_{t=1}^{T}\log\pi_{\theta}(a_{t}\mid\mathbf{a}_{1:t-1})\right]$, which is the sum of log probabilities, whose expected value (i.e., the average log probability) with respect to $\mathbf{\theta}$ should be equal to 0, because the probability distribution $\pi_{\mathbf{\theta}}$ is fixed under its own expectation. The proof is thus completed. ∎
Since the baseline value introduced by the QFR algorithm is determined by the reward of the greedy policy, and corresponds to the reward distribution, it is statistically independent of the samples sampled by the normal policy, satisfying the requirements for the proof of unbiasedness.
@@
The transition model of MDPs defined in Section III-B can be decomposed as follows. Firstly, the policy model gives a token $\mathbf{a}_{t}$ based on the input sequence $\mathbf{s}_{t}=\mathbf{a}_{1:t-1}$, and the output sequence $\mathbf{s}_{t+1}^{\prime}$ is built by simply appending $\mathbf{a}_{t}$ to $\mathbf{s}_{t}$, i.e., $\mathbf{s}_{t+1}^{\prime}=\mathbf{a}_{1:t}$. If the transition is deterministic, meaning the transition function satisfies the Dirac distribution, then $\mathbf{s}^{\prime}_{t+1}$ is directly used as the input state for next step. If a probabilistic transition function is used, $\mathcal{T}(\mathbf{s}^{\prime\prime}_{t+1}|\mathbf{s}_{t},a_{t})$ is applied and it generates a new random sequence $\mathbf{s}^{\prime\prime}_{t+1}$ as the next input state.
@@
This proposition demonstrates that MDPs corresponding to state transfer functions that satisfy the Dirac distribution have the smallest variance compared to MDPs with diverse state transfer functions, and thus REINFORCE exhibits an advantage in deterministic environments. The lower environmental variance alleviates the deficiency of high variance in the REINFORCE algorithm.

### v2

@@ III-D Comparing REINFORCE with PPO
$=\mathbb{E}_{\tau\sim p_{\mathbf{\theta}}(\tau)}\left[\sum_{t=0}^{T}\nabla_{\mathbf{\theta}}\log\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)R(\tau)\right]$
$=\mathbb{E}_{\mathbf{a}_{1:t}\sim\pi_{\mathbf{\theta}}}\left[\sum_{t=0}^{T}s_{\mathbf{\theta}}\left(\mathbf{a}_{1:t}\right)r\left(\mathbf{a}_{1:T}\right)\right],$ (6)
@@
where $\tau$ represents a trajectory, $p_{\mathbf{\theta}}(\tau)$ represents the probability of trajectory $\tau$ sampled by policy $\pi_{\mathbf{\theta}}$. In the above MDP, the current policy is $\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$, and thus we have $p_{\mathbf{\theta}}(\tau)=p\left(a_{0}\right)\prod_{t=0}^{T}\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$. $R(\tau)$ represents the cumulative reward of trajectory $\tau$. Since IC can only be calculated after the complete expression is generated, only the reward for the final step is non-zero. Therefore, $R(\tau)=r\left(\mathbf{a}_{1:T}\right)$. Let the score function be denoted as $s_{\mathbf{\theta}}\left(\mathbf{a}_{1:t}\right)=\nabla_{\mathbf{\theta}}\log\pi_{\mathbf{\theta}}\left(a_{t}\mid\mathbf{a}_{1:t-1}\right)$.
@@
Here $r\left(\mathbf{a}_{1:t}\right)$ is used as the metric of the effectiveness of the current policy. It can also be replaced by the following five metrics: $\sum_{t=0}^{\infty}r_{t}$ (total reward of the trajectory)[52], $\sum_{t^{\prime}=t}^{\infty}r_{t^{\prime}}$ (reward following action $a_{t}$)[52], $\sum_{t^{\prime}=t}^{\infty}r_{t^{\prime}}-b(\mathbf{s}_{t})$ (baselined version of the previous formula)[52], $Q^{\pi}(\mathbf{s}_{t},a_{t})$ (state-action value function)[53], $r_{t}+V^{\pi}(\mathbf{s}_{t+1})-V^{\pi}(\mathbf{s}_{t})$ (TD residual)[54], and $A\left(\mathbf{s}_{t},a_{t}\right)=Q^{\pi}(\mathbf{s}_{t},a_{t})-V^{\pi}(\mathbf{s}_{t})$ (advantage function)[54]. PPO [21] adopts the advantage function to measure the effectiveness of the policy under the actor-critic framework of TD[55], and was employed in [19]. Both the actor network (the policy) and a critic network (the value network) are trained at the same time. Generally, the merits of TD methods over the Monte Carlo algorithms[56] to update network weights is that data from each step can be used for training, increasing the sample efficiency.
Unfortunately, the MDP considered in this work is with only trajectory feedback, i.e., non-zero rewards are only received at the final step. This means that the advantage function, as well as PPO, becomes less effective, since the critic network is hard to train and involves high biases. Comparatively, Monte Carlo algorithms without critic networks such as REINFORCE are better suited to. Generally, the REINFORCE calculates the policy gradient based on (6) using the Monte Carlo method:
@@
where $a_{t}^{i}\sim\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{a}_{1:t-1}^{i}\right)$ for all $i=1,\ldots,N$. It is essentially a likelihood maximization weighted by rewards using samples extracted from rollouts. And for $r\left(\mathbf{a}_{1:t}\right)$, REINFORCE normally adopts one of the first three metrics aforementioned and discards the critic network, thus is an unbiasesd estimation of policy gradient. We will delve into the theoretic details in Section IV-C.
IV Steadier RL for Alpha Mining
In light of the potentially high biases caused by PPO, this section introduces a novel REINFORCE-based method for formulaic alpha factor mining, termed QuantFactor REINFORCE (QFR). Due to the Monte Carlo nature, QFR estimates unbiased policy gradient. Considering the high variance of REINFORCE methods, an effective greedy baseline is proposed, theoretically grounded by an upper bound on the variance. We demonstrate that QFR exhibits reduced variance compared to REINFORCE. The proof that QFR, when applied to MDPs with deterministic transition function, exhibits the lowest variance compared to those with probabilistic transition function, is also provided. These theoretical analysis supports a steady mining process of formulaic alpha factors. Furthermore, the Information Ratio (IR) is introduced as a new reward shaping mechanism to further promote the steadiness.
Figure 2: Comparison between QFR and PPO. By discarding the critic network (value model), QFR requires much fewer rollouts than PPO, which significantly improves the overall convergence speed.
@@ IV-A The Proposed Algorithm
Inspired by the REINFORCE algorithm with baseline[24], the proposed QFR optimizes the gradient estimation by introducing a specific baseline value to reduce the variance:
@@
where the baseline value $r\left(\mathbf{\bar{a}}_{1:T}^{i}\right)$ can be obtained by greedily sampling a formulaic alpha factor with the maximal conditional probability and calculating the associated reward, i.e., $\bar{a}_{t}\in\operatorname{argmax}\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{\bar{a}}_{1:t-1}\right)$. The proposed baseline serves as a normalizing function by comparing the reward of the random response with that of the greedy response, thus can reduce the variance in the gradient estimation. While PPO can also reduce large variance during training, it does so at the cost of training an additional value model with similar complexity to the policy model. This not only slows down the convergence, but substantially increases the sampling time costs, as a forward pass must be performed for each state sampled from the buffer to obtain $V^{\pi}(\mathbf{s}_{t})$. Comparatively, QFR receives the rewards as a trajectory feedback, which fits the nature of the underlying MDP, and saves much computational costs for rollout. The differences between QFR and PPO based formulaic alpha factors mining frameworks are depicted in Figure. 2
The detailed pipeline of the QFR algorithm is shown in Fig. 3. The policy model generates one new token (action) at each step, and the sequence of generated tokens (states) can be uniquely converted into an expression based on RPN. The policy model uses random sampling and greedy sampling to complete two trajectories and generate two expressions, i.e., two factors. The combination model is used to maintain a weighted combination of principal factors and, at the same time, to evaluate these two factors. Specifically, the combined factor values are used to calculate $r\left(\mathbf{a}_{1:T}\right)$ and $r\left(\mathbf{\bar{a}}_{1:T}\right)$ along with the real asset data features $\mathcal{X}$. These two rewards, calculated by the proposed shaped reward function (detailed in Section IV-B), form the estimation of (8). The algorithmic steps are given in Algorithm 1.
Input: Real asset price dataset $\mathcal{Y}$, Real asset feature dataset $\mathcal{X}=\{\mathbf{X}^{\prime}_{l}\}$, Initial policy weight $\mathbf{\theta}$, Initial combination model weight $\mathbf{\omega}$, Reward balancing coefficient $\lambda$.
@@
1while not converged do
2       Construct a normal factor $f_{n}$ with $\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{a}_{1:t-1}\right)$;
3       Construct a greedy factor $f_{g}$ with $\pi_{\mathbf{\theta}}\left(\cdot\mid\mathbf{\bar{a}}_{1:t-1}\right)$;
4       Compute normal factor values $\{\mathbf{z}_{n,l}\}=\{f_{n}(\mathbf{X}_{l})\}$;
5       Compute greedy factor values $\{\mathbf{z}_{g,l}\}=\{f_{g}(\mathbf{X}_{l})\}$;
6       Compute $\{\mathbf{z}^{\prime}_{n,l}\}$ and $\{\mathbf{z}^{\prime}_{g,l}\}$ of the normal factor and greedy factor with $\mathbf{\omega}$, respectively;
7       Compute $r\left(\mathbf{a}_{1:T}\right)$ and $r\left(\mathbf{\bar{a}}_{1:T}\right)$ with the shaped reward and both $\{\mathbf{z}^{\prime}_{n,l}\}$ and $\{\mathbf{z}^{\prime}_{g,l}\}$ via Eq. (11);
@@ 5       Compute greedy factor values $\{\mathbf{z}_{g,l}\}=\{f_{g}(\mathbf{X}_{l})\}$;
To our best knowledge, we are the first to introduce this baseline in MDPs for formulaic alpha factors mining. Notably, while this baseline looks straightforward, it works very well for the underlying MDP. Moreover, we provide comprehensive and detailed theoretical understandings of why this baseline is particularly well suited to these MDPs in Section IV-C.
Figure 3: The detailed pipeline for the proposed QFR algorithm.
Figure 4: Clip function values versus training time steps, with an example of $\overline{IR}$ values. When the clipped function value, i.e., the $\overline{IR}$ test value, is greater than the $\overline{IR}$ value of the generated combined factor, the factor quality is assumed to be poor, and the last step of the trajectory will receive a discounted $\overline{IC}$ as a reward. Conversely, a complete $\overline{IC}$ is obtained as a reward.
IV-B Time-varying Reward Shaping
The reward function in [19] focuses on the absolute excess returns of factors, while ignoring the risk-adjusted characteristics of the returns, i.e., how the factors are resistant to the volatility of the market. In order to better balance returns and risks, our work not only evaluate the predictive accuracy of factors, but considers the stability of their predictive signals, thus providing a more comprehensive assessment of factor performance.
Specifically, we introduce the Information Ratio (IR) for reward shaping. IR is a financial metric that measures the excess returns of factors relative to excess risk and is commonly used to evaluate the risk-adjusted performance of the factors:
@@ IV-B Time-varying Reward Shaping
Equation (9) describes how the IR of a single factor is calculated. After the combination model, the IR for the factors becomes:
@@
Equation (11) is the proposed shaped reward. $\mathbb{I}$ denotes the indicator function, which takes the value of $1$ if the condition inside the parentheses is satisfied, and 00 otherwise. $\alpha$ is the time delay, set to $9\times 10^{4}$ time steps, $\eta$ is the slope of the change in the IR test value, set to $2.65\times 10^{-6}$, $\delta$ is the maximum IR test value, set to 0.3, and $\lambda$ is used to balance the main reward with the shaped reward, set to 0.02. The varying of the clip function over training time is illustrated in Fig. 4.
Using the IR test as reward shaping can force the QFR algorithm focusing more on the long-term stability and consistency of the factors during optimization, rather than just the short-term profit peaks. This reward shaping method guides the policy network to find factors that maintain stable predictive capability across different market conditions, reducing the risk of overfitting the training data. It helps to reduce the model’s over-reliance on specific market conditions, thereby enhancing the generalization ability of factors across different market environments.
@@ IV-C The Theoretical Analysis
We provide a set of theoretical results for Algorithm 1, which includes the analysis of the training variance under different state transition functions in Proposition 2, the derivation of the upper bound of the variance in Proposition 3, and the demonstration that the variance decreases relative to the REINFORCE algorithm in Proposition 4. To justify the algorithm design, we first prove that the variance of the proposed algorithm is bounded.
@@
where (12) follows the so-called log-derivative trick, (13) follows the idea of Bartlett identity $\sum_{z}\nabla_{\mathbf{\theta}}p_{\mathbf{\theta}}(z)b=\nabla_{\mathbf{\theta}}\left[\sum_{z}p_{\mathbf{\theta}}(z)b\right]=\nabla_{\mathbf{\theta}}[1\cdot b]=0$, in which $p_{\mathbf{\theta}}$ can be any distribution and $b$ is constant[57]. Notice that $\mathbf{\bar{a}}_{1:T}$ is conditionally independent on $\mathbf{\theta}$, due to the greedy sampling and apply $p_{\mathbf{\theta}}$ to the the distribution $\pi_{\mathbf{\theta}}\left(\bar{a}_{1:T}\mid x\right)$. Then, regarding the second item, we only need to consider $\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\theta}}\left[\nabla_{\theta}\sum_{t=1}^{T}\log\pi_{\theta}(a_{t}\mid\mathbf{a}_{1:t-1})\right]=\nabla_{\theta}\mathbb{E}_{\mathbf{a}_{1:T}\sim\pi_{\theta}}\left[\sum_{t=1}^{T}\log\pi_{\theta}(a_{t}\mid\mathbf{a}_{1:t-1})\right]$, which is the sum of log probabilities, whose expected value (i.e., the average log probability) with respect to $\mathbf{\theta}$ should be equal to 0, because the probability distribution $\pi_{\mathbf{\theta}}$ is fixed under its own expectation. The proof is thus completed. ∎
Since the baseline value introduced by QFR is determined by the reward of the greedy policy, and corresponds to the reward distribution, it is statistically independent of the samples sampled by the normal policy, satisfying the requirements for the proof of unbiasedness.
@@
The transition model of MDPs defined in Section III-C can be decomposed as follows. Firstly, the policy model gives a token $\mathbf{a}_{t}$ based on the input sequence $\mathbf{s}_{t}=\mathbf{a}_{1:t-1}$, and the output sequence $\mathbf{s}_{t+1}^{\prime}$ is built by simply appending $\mathbf{a}_{t}$ to $\mathbf{s}_{t}$, i.e., $\mathbf{s}_{t+1}^{\prime}=\mathbf{a}_{1:t}$. If the transition is deterministic, meaning the transition function satisfies the Dirac distribution, then $\mathbf{s}^{\prime}_{t+1}$ is directly used as the input state for next step. If a probabilistic transition function is used, $\mathcal{T}(\mathbf{s}^{\prime\prime}_{t+1}|\mathbf{s}_{t},a_{t})$ is applied and it generates a new random sequence $\mathbf{s}^{\prime\prime}_{t+1}$ as the next input state.
@@
This proposition shows that MDPs with deterministic state transition functions (following the Dirac distribution) exhibit lower variance compared to those with probabilistic state transition functions. Consequently, this helps mitigate the high variance issue inherent in REINFORCE.